How would I do this kind of problem
A telephone number consists of seven digits, the first three representing the exchange. How many different telephone numbers are possible within the 537 exchange?

Answer by Nate(3500)

Depends, can you repeat digits?
If you can ....
5 - 3 - 7
Three numbers:
3 * 3 * 3 = 27 different possibilites
3 * 3 * 3 * 9 * 9 * 9 * 9 = 177,147
If you can not ....
3 * 2 * 1 = 6 different possibilites
3 * 2 * 1 * 7 * 6 * 5 * 4 = 5,040